• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Hall A (Techs)

    Log entry time 16:52:49 on November 13, 2014

    Entry number 403976


    "Checklist complete and target window and spectrometer guards are removed"

    "Angle limits for the left spectrometer are 23 degrees to 23 degrees"

    "Angle limits for the right spectrometer are 67.25 degrees to 67.25 degrees"

    "The tech on call is JESSIE BUTLER"

    "Sieve Slit is NOT installed on right/left spectrometers"

    "Right spectrometer window and VDC covers are NOT removed, see Jack for more information"

    "Right spectrometer magnets are off per jack"

    "Left collimator is in the "open" position"

    "The left Dipole is having software problems, Jack has been notified, it has not been ramped"